# On the Fairness, Diversity and Reliability of Text-to-Image Generative Models

### v1

@@ 4 Results
Reliability, fairness and diversity are key T2I model traits. We evaluate both intentionally-biased and presumed-benign models, marking an improvement over related works [48, 17]. Our method quantifies ethical AI metrics, which we exploit for detecting the presence and provenance of intentional biases. Recognizing that bias triggers may never appear in user prompts, we view intentional, extreme bias as a result of intentionally unfair or deterministic training practices, which reflects in generated outputs.
Figure 9: Qualitative impact of our leave-one-out experiments which underpin our generative fairness $\mathcal{F}_{\tilde{x}_{T}}$ evaluations. Quantitative results report the low guidance configuration. Here we show both. The top row of each example is the base model [11] output.
@@
Ablation: Generation Steps vs. Reliability. As discussed prior, diffusion-based models remove noise over a period ‘$T$’, with the quality of image generally increasing $\propto T$. To ensure that our global and local reliability evaluations are not adversely affected by increased generation steps, we conduct an additional ablation study for SD-V1.5, BAGM and TPA models to highlight the relationship between $\mathcal{R}_{G}$, $\mathcal{R}_{L}$ and the generation time $T$. We double the generation time (50$\rightarrow$100 steps), maintaining the same inference time variables used prior. We visualize $\mathcal{P}_{G,L}(\varphi_{s})$ comparisons in Fig. 6. The relationship between presumed-benign and intentionally-biased models (for the same $T$) is relatively consistent as per Fig. 6. From this ablation study we can infer that semantic perturbations behave consistently and intentionally-biased models display consistent behaviors w.r.t. base models at different denoising steps.
@@
Model Type $\varphi_{Mo}$ $\mathcal{P}_{G}(\varphi_{s}=\varphi_{Mo})$ $\varphi_{Mo}$ $\mathcal{P}_{L}(\varphi_{s}=\varphi_{Mo})$
@@
Ablation: Vision Transformer (ViT) Comparison. For our main evaluations, we use the CLIP Vision Transformer (ViT)-B/32 [29] for image similarity calculations. The ViT extracts and projects features from the pixel domain to the ViT domain, where we compute the cosine similarity of image pairs. High cosine similarity indicates aligned images, which we can visualize through our $\mathcal{D}_{\tilde{x}_{T}}$ image similarity matrices (see Figs. 7 and 8). To assess the impact of the deployed ViT, we performed ablation studies with two additional models: Google ViT (GViT) [49] and Facebook DeiT (FBViT) [44]. We repeated our global and local reliability experiments with a smaller test prompt set to confirm that our evaluations are not ViT-dependent.
@@
The construction of each ViT feature space is different and thus, the $\varphi_{s}$ required to change the image will also differ across models. However, the relationship between benign vs. intentionally-biased models should remain consistent. We report ViT-dependent $\mathcal{P}_{G,L}$ Modal characteristics in Table 3 and observe that while the semantic perturbations required for each model does differ (as expected), the relationship between presumed-benign and intentionally-biased models is consistent, demonstrating ViT independence. Analyzing Table 3, the CLIP ViT model requires stronger semantic perturbations to induce significant image changes in $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ evaluations, unlike the Google ViT and Facebook DeiT models, which require less perturbations. We hypothesize that the higher $\varphi_{s}$ in CLIP ViT reflects a broader representational space compared to GViT and FBViT. This suggests that CLIP ViT-B/32 [29] has a wider space, which makes it well-suited for comparing image pairs after applying semantic perturbations.
@@ 5 Conclusion
Most models are susceptible to unreliability under certain input conditions, with intentionally-biased and backdoored models often especially unreliable and unfair. This work demonstrates how semantic perturbations can reveal such behavior in text-to-image models, providing a quantitative measure of global and local reliability. Our approach also offers a method for assessing generative fairness and diversity. Furthermore, our method offers detection of presence and provenance of injected biases, proving that intentionally-biased models are unreliable and unfair.
@@ 6 Acknowledgments
This research and Dr. Jordan Vice are supported by the NISDRG project #20100007, funded by the Australian Government. Dr. Naveed Akhtar is a recipient of the ARC Discovery Early Career Researcher Award (project #DE230101058), funded by the Australian Government. Professor Ajmal Mian is the recipient of an ARC Future Fellowship Award (project #FT210100268) funded by the Australian Government.

### v2

@@ 4 Results
We evaluate both intentionally-biased and presumed-benign models, marking an improvement over related works [56, 57]. Our method quantifies ethical AI metrics, which we exploit for detecting the presence and provenance of intentional biases. Recognizing that bias triggers may never appear in user prompts, we view intentional, extreme bias as a result of intentionally unfair or deterministic training practices, which reflects in generated outputs.
4.1 Ablation Studies
@@
Generation Steps vs. Reliability. As discussed prior, diffusion-based models remove noise over a period ‘$T$’, with the quality of image generally increasing $\propto T$. To ensure that our global and local reliability evaluations are not adversely affected by increased generation steps, we conduct an additional ablation study for SD-V1.5, BAGM and TPA models to highlight the relationship between $\mathcal{R}_{G}$, $\mathcal{R}_{L}$ and the generation time $T$. We double the generation time (50$\rightarrow$100 steps), maintaining the same inference time variables used prior. We visualize $\mathcal{P}_{G,L}(\varphi_{E})$ comparisons in Fig. 6. The relationship between presumed-benign and intentionally-biased models (for the same $T$) is relatively consistent as per Fig. 6. From this ablation study we can infer that embedding perturbations behave consistently and intentionally-biased models display consistent behaviors w.r.t. base models at different denoising steps.
Input Prompt Distribution vs. Reliability Given that conditional image generation is largely governed by the input prompt, we consider evaluating our reliability method using the Google Conceptual Captions (GCC) dataset [64] for the baseline SD 1.4/5/2.1 models. Through this evaluation, we aim to determine if the comparisons across models are consistent with our original findings in Table 2. We report our GCC probability distribution characteristics in Table 4. Comparing results across the two tables, we observe that the global reliability ‘$\mathcal{R}_{G}$’ follows a similar trend, regardless of input distribution. Through iterative developments, models are becoming more reliable on a global scale, regardless of input prompt. SD-V2.1 (the latest of the three) typically requires more globally-applied embedding perturbations to cause a significant change in the output. In comparison, the difference in $\mathcal{R}_{L}$ may demonstrate that the input distribution may have an impact on local reliability. Logically, this is supported by the construction of prompts and the corresponding semantics within a sentence.
@@
Model Type $\varphi_{Mo}$ $\mathcal{P}_{G}(\varphi_{E}=\varphi_{Mo})$ $\varphi_{Mo}$ $\mathcal{P}_{L}(\varphi_{E}=\varphi_{Mo})$
SD-V1.4 benign 0.1158 5.9330 0.3011 2.4991
SD-V1.5 benign 0.1178 6.4769 0.2722 2.5792
SD-V2.1 benign 0.1368 6.2498 0.4020 2.2244
Table 4: Global and local model reliability ‘$\mathcal{R}_{G/L}$$\varphi_{E}$-distribution comparison when generating images over $T=50$ denoising steps, using prompts from the Google Conceptual Captions (GCC) [64] dataset. ‘$\varphi_{Mo}$’ and ‘$\mathcal{P}_{G/L}(\varphi_{E}=\varphi_{Mo})$’ describe the Modal value and Mode, respectively. We visualize these distributions in Fig. 6.
$\mathcal{R}_{G}$ $\mathcal{R}_{L}$
Model Type $\varphi_{Mo}$ $\mathcal{P}_{G}(\varphi_{E}=\varphi_{Mo})$ $\varphi_{Mo}$ $\mathcal{P}_{L}(\varphi_{E}=\varphi_{Mo})$
@@
Figure 9: Comparison of probability distribution functions $\mathcal{P}_{G/L}(\cdot)$ for benign vs. BAGM [16] models when using three unique ViTs to demonstrate that choice of ViT does not adversely impact our evaluations. Reported Modal characteristics in 2 are inferred from the curves represented here, where: (a) visualizes global reliability curves and, (b) visualizes local reliability curves.
@@
Vision Transformer (ViT) Comparison. For our main evaluations, we use the CLIP Vision Transformer (ViT)-B/32 [38] for image similarity calculations. The ViT extracts and projects features from the pixel domain to the ViT domain, where we compute the cosine similarity of image pairs. High cosine similarity indicates aligned images (see Fig. 7). To assess the impact of the deployed ViT, we performed ablation studies with two additional models: Google ViT (GViT) [65] and Facebook DeiT (FBViT) [66]. We repeated our global and local reliability experiments with a smaller test prompt set to confirm that our evaluations are not ViT-dependent.
@@
The construction of each ViT feature space is different and thus, the $\varphi_{E}$ required to change the image will also differ across models. However, the relationship between benign vs. intentionally-biased models should remain consistent. We report ViT-dependent $\mathcal{P}_{G,L}$ Modal characteristics in Table 5 and observe that while embedding perturbations required for each model does differ (as expected), the relationship between presumed-benign and intentionally-biased models is consistent, demonstrating ViT independence. Analyzing Table 5, the CLIP ViT model requires stronger perturbations to induce significant image changes in $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ evaluations, unlike the Google ViT and Facebook DeiT models, which require less perturbations. We visualize the corresponding probability distribution functions in Fig. 9 and 10. We hypothesize that the higher $\varphi_{E}$ in CLIP ViT reflects a broader representational space compared to GViT and FBViT. This suggests that CLIP ViT-B/32 [38] has a wider space, which makes it well-suited for comparing image pairs after applying embedding perturbations.
Figure 10: Comparison of probability distribution functions $\mathcal{P}_{G/L}(\cdot)$ for benign vs. TPA [1] models when using three unique ViTs to demonstrate that choice of ViT does not adversely impact our evaluations. Reported Modal characteristics in 2 are inferred from the curves represented here, where: (a) visualizes global reliability curves and, (b) visualizes local reliability curves.
Generative Diversity vs. Abstraction Level. Ontology-based hierarchies have been used to structure semantic concepts in image understanding tasks [67, 68]. We adopt a similar approach here to elaborate on our generative diversity metric. Concept representations and the diversity of training data will manifest in the output space of generative models. In cases where the level of abstraction of a concept is high (like with high-level classes or single character tokens), we expect that the diversity should be high. In comparison, high specificity concepts like animal breeds will logically constrain the output space and result in low diversity images. Thus, understanding the specificity and ontology of $\tilde{x}_{T}$ is important for providing additional context to the diversity evaluation.
We visualize this phenomenon in Figs. 11 and 12 for animal and food ontologies, respectively. We observe that as the level of abstraction decreases from level $i\rightarrow k$, so too does $\mathcal{D}_{\tilde{x}_{T}}$. The higher the specificity, the lower the diversity as shown when we compare the concept of animal to “polar bear” in Fig. 11, which shows a reduction in $\mathcal{D}_{\tilde{x}_{T}}$ of 0.287 points. This supports our logic for identifying intentional biases as visualized previously (see Fig. 7). Similarly in Fig. 12, we again observe that an inverse relationship exists between $\mathcal{D}_{\tilde{x}_{T}}$ and the level of abstraction. As expected, we see that the class ‘carrot’ is less diverse than the parent classes vegetable and food. This study is also crucial for understanding cases where for example “elephant” and “toilet” demonstrate an expected, low $\mathcal{D}_{\tilde{x}_{T}}$ as previously reported in Table 3. Lower $\mathcal{D}_{\tilde{x}_{T}}$ score for a concept across models is indicative of a decrease in abstraction and a bias toward a particular representation. This is a common goal of bias injection methodologies and our diversity metric can be used to find uncharacteristically homogeneous representations, as is the case for BAGM and TPA methods reducing the diversity of bias trigger concepts by $0.068$ and $0.288$ points, respectively (see Fig. 7).
Figure 11: Comparison of the level of abstraction vs. generative diversity across the animal ontology. As the abstraction level decreases from $i\rightarrow k$, so too (logically) does the diversity. This is attributed to the increase in specificity. Images are generated using the benign SD-V2.1 model.
Figure 12: Comparison of the level of abstraction vs. generative diversity across the food ontology. Images are generated using the benign SD-V2.1 model. Similar to Fig. 11, we see that traversing down the tree logically increases the specificity of the concept, at the cost of generated image diversity.
In Table 6, we report how sub-classes ‘$C_{i,j}$’ demonstrate lower diversity than their parent class $C_{i}$. We observe for all pairs, the sub-class has less diverse outputs than the parent class. This behavior is expected for general classes. We recall that for the intentionally-biased SD-V1.5 (BAGM) model, the bias trigger ‘drink’ reported $\mathcal{D}_{\tilde{x}_{T}}=0.150$ and ‘coffee’ reported $\mathcal{D}_{\tilde{x}_{T}}=0.226$ (as reported in Table 3). Comparing these results to the relevant class/sub-class pairs in Table 6, we see that intentional biases cause a decrease in abstraction to a point where the output space is clearly constrained as it pertains to the sensitive, bias trigger tokens.
While our diversity metrics target bias provenance, this ontological ablation reveals a secondary use i.e., understanding how generative models organize semantic concepts and how diverse output representations are in general.
$C_{i}$ $\mathcal{D}_{C_{i}}$ $C_{i,j}$ $\mathcal{D}_{C_{i,j}}$ $C_{i}$ $\mathcal{D}_{C_{i}}$ $C_{i,j}$ $\mathcal{D}_{C_{i,j}}$
ape 0.2069 silverback gorilla 0.0597 broccoli 0.1210 steamed broccoli 0.0931
bear 0.1804 polar bear 0.0617 strawberry 0.1327 strawberry jam 0.1047
bird 0.2255 bald eagle 0.0836 carrot 0.1021 roasted carrot 0.0825
fish 0.2734 trout 0.1408 banana 0.1978 banana smoothie 0.0899
whale 0.1172 orca 0.0649 coffee 0.3308 starbucks 0.1943
shark 0.1896 hammerhead shark 0.1040 soda 0.3758 coca cola 0.2830
Table 6: Diversity characteristics when progressing through class hierarchies from the parent class $C_{i}$ to the sub-class ‘$C_{i,j}$’. As shown, traversing down the ontology tree and increasing specificity yields lower diversity output representations, as expected. The SD-V2.1 model was used to generate images for these evaluations.
5 Limitations
Comprehensive reliability evaluations of any system requires some form of stress test and an increase in the load. Applying a series of embedding perturbations comes at a cost of increased computational complexity. By increasing the number of generated images per step and reducing the perturbation step size, the evaluations can become more robust, but at the cost of time. In comparison, increasing the perturbation step size and/or reducing the number of images will result in less comprehensive reliability evaluations and a lower-resolution understanding of the learned embedding space. This presents an interesting optimization problem. Furthermore, our global and local reliability evaluations do assume grey-box conditions, meaning that they require access to the text-encoder output. However, generative fairness and diversity evaluations are not constrained by these assumptions and could be leveraged for assessing black-box generative models as well. With that in mind, our fairness and diversity evaluations and metrics can be extended to any input prompts or concepts - not only those that exhibit unreliable model behavior. From a definition perspective, Fairness and Diversity are typically associated with social biases and ethical AI implementations. While not necessarily a limitation, it is worth noting that our definitions of each consider a generalized approach. We instead define both w.r.t. the influence that concepts have on conditioning and thus, output representations
@@ 6 Conclusion
Increased popularity in generative models exposes fairness and reliability concerns. With models gaining wide public attention, it is important to consider their fairness, diversity and reliability. Most models are susceptible to unreliability under certain input conditions, with intentionally-biased and backdoored models often especially unreliable and unfair. This work demonstrates how embedding perturbations can reveal such behavior in text-to-image models, providing a quantitative measure of global and local reliability. Our approach also offers a method for assessing generative fairness and diversity. Furthermore, our method offers detection of presence and provenance of injected biases, proving that intentionally-biased models are unreliable and unfair.
@@ 7 Acknowledgments
This research was supported by National Intelligence and Security Discovery Research Grants (project #NS220100007), funded by the Department of Defence, Australia. Dr. Naveed Akhtar is a recipient of the Australian Research Council Discovery Early Career Researcher Award (project number DE230101058) funded by the Australian Government. Professor Ajmal Mian is the recipient of an Australian Research Council Future Fellowship Award (project number FT210100268) funded by the Australian Government.